A chemist combines HCl acid solutions of 40%, 50%, and 25% to get 200 liters of a 32% HCl solution. If the chemist uses twice as much of the 25% HCl solution as of the 40% HCl solution, find how many liters of each kind he uses.
Answer by chessace(471)

Translate:
a+b+c=200
.4a+.5b+.25c=.32*200
c=2a
Compute:
b=200-3a
.4a+.5(200-3a)+.25*2a=.32*200
(.4-1.5+.5)a=200(.32-.5)
-.6a=200*(-.18)
a=200*.3= 60 liters
c= 120 liters
b= 20 liters